# Convergence analysis of Levenberg-Marquardt method with Singular Scaling for non-zero residue nonlinear least-squares problems [^1] This work is supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360]

### v1

@@ 3 Auxiliary Results
Now let us show that $d_{R}\notin\mathcal{N}(L)$. Suppose $d_{R}\in\mathcal{N}(L)$, from (26), we get $\left\lVert Ld_{N}\right\rVert^{2}=-d_{N}^{T}L^{T}Ld_{R}=0,$ implying that $d_{N}\in\mathcal{N}(L)$ and $d_{N}\in\mathcal{N}(J_{k})$, which is a contradiction with Assumption 1.
By Assumption 1, there exists $\gamma$ such that $\gamma\left\lVert d_{N}\right\rVert^{2}\leq\left\lVert J_{k}d_{N}\right\rVert^{2}+\left\lVert Ld_{N}\right\rVert^{2}$, which implies
@@
$\left\lVert d_{k}\right\rVert^{2}$ $=\left\lVert d_{N}+d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+2d_{N}^{T}d_{R}+\left\lVert d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}$ (32)
@@
$\leq\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}\right\rVert^{2}=\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}-J_{*}^{T}F_{*}\right\rVert^{2}$
@@
Since $J(x)$ is continuous, it is clear that for a sufficiently small $\delta>0$, condition (13) is satisfied for all $x_{k}\in B(x^{*},\delta)$. Finally, we have
@@
Lemma 6.
Suppose that Assumptions 13 are valid in $B(x^{*},\delta)$, for some $\delta>0$. If $x_{k}\in B(x^{*},\delta)$, $rank(J_{k})=\ell\geq rank(J_{*})=q\geq 1$, and Assumption 4 is satisfied, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r},$ with $r\in(0,1]$, then there exists $c_{1}>0$ such that $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$.
Proof.
Consider the GSVD for the pair $(J_{k},L)$. In this case where $rank(J_{k})=\ell,rank(J_{*})=q\geq 1$ and $rank(L)=p$ with $n-q\leq p$ and $n-\ell\leq p$ (see Remark 1), we remark that:
@@
  • For $n-q\leq i\leq p$, then $(\Gamma_{k})_{ii}=\frac{1}{\sigma_{i,k}+\lambda_{k}\mu_{i,k}}<\frac{1}{\sigma_{i,k}}\leq\frac{1}{\sigma^{*}_{\text{min}}},$ where $\sigma^{*}_{\text{min}}$ is the smallest positive singular value of $J(x^{*}).$
Note that, for $x_{k}\in B(x^{*},\delta)$ and due to the continuity of $\mu_{i,k}$, in the case where $n-\ell\leq i<n-q$ we can assume that $\mu_{i,k}>\frac{1}{2}$ and thus, for $\delta$ sufficiently small, we have $(\Gamma_{k})_{ii}\leq\frac{1}{4\lambda_{k}}$. Therefore,
$\left\lVert\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]\right\rVert\leq\max{\Big{[}\frac{1}{\lambda_{k}},\frac{1}{4\lambda_{k}},\frac{1}{\sigma^{*}_{\text{min}}},1\Big{]}}\leq\frac{1}{\lambda_{k}},$ (33)
as $x_{k}\in B(x^{*},\delta)$ and $\left\lVert J_{k}^{T}F_{k}\right\rVert<1.$
From (19) and (33), we obtain
@@
$=\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}\Big{(}{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})\Big{)}\right\rVert$
$\leq\left\lVert X_{k}\right\rVert^{2}\frac{1}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$
@@
$\left\lVert d_{k}\right\rVert$ $\leq\left\lVert X_{k}\right\rVert^{2}\frac{1}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$ (36)
@@
$\leq\frac{1}{\gamma\lambda_{k}}L_{4}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{2}+\frac{1}{\gamma\lambda_{k}}\left\lVert J_{k}^{T}F(\bar{x}_{k})\right\rVert+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
Finally, from Assumption 4 and $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, the result follows. ∎
@@ 4 Local Convergence
Lemmas 5 and 6 showed that $\left\lVert d_{k}\right\rVert\leq c_{1}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$, under specific assumptions. As we shall see in this section, such inequality is key for the local convergence analysis under an error bound condition (Assumption 3).
@@
Throughout this section, we consider the LMMSS iteration (2)–(3) with $\alpha_{k}=1$ for every $k$, that is, the “pure” LMMSS.
Lemma 7.
Let $\{x_{k}\}$ be the sequence generated by the LMMSS (with $\alpha_{k}=1$) and suppose Assumptions 1-3 hold. If $x_{k+1},x_{k}\in B(x^{*},\delta)$ and $\left\lVert d_{k}\right\rVert\leq c_{1}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$, then
@@
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq L_{4}\|d_{k}\|^{2}+L_{0}L_{2}\|x_{k}-\bar{x}_{k}\|\|x_{k+1}-\bar{x}_{k}\|+\|J_{k}^{T}F(\bar{x}_{k})\|$ (42)
@@ 4.1 Constant Rank
In this section, we assume that $1\leq\text{rank}(J(x^{*}))=q\leq\min\{n,m\}$, and $\text{rank}(J(x))=\text{rank}(J(x^{*}))$ for each $x\in B(x^{*},\delta)$. From Lemma 5, we recall that $\left\|d_{k}\right\|\leq c_{1}\text{dist}(x_{k},X^{*})$, with $c_{1}=\Big{(}\sqrt{\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1}\Big{)}\frac{\kappa}{\lambda_{p}^{*}}L_{3}$, provided that $\lambda_{k}>0,$ for $x_{k}\neq x^{*}$.
@@
Lemma 8.
Suppose that Assumptions 1-3 are valid, and $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$. Additionally, assume that Assumption 4 is verified, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|$. If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, with
$\varepsilon\leq\left(\frac{\eta\omega}{\tilde{C}}\right)^{1/r},$ (43)
with $\eta\in(0,1)$ and $L_{6}=L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1}$, then
@@
Proof.
We know that $J(x)$ has constant rank in $B(x^{*},\delta)$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, and Assumptions 4 is satisfied. Thus, from (37) and (7), we have
@@
$\textup{dist}(x_{k+1},X^{*})\leq\eta\textup{dist}(x_{k},X^{*}).$
Lemma 9.
@@
Proof.
The proof is done by induction. For $k=0$, we have $\textup{dist}(x_{0},X^{*})\leq\|x_{0}-x^{*}\|\leq\varepsilon$ and
$\|x_{1}-x^{*}\|\leq\|x_{1}-x_{0}\|+\|x_{0}-x^{*}\|\leq\|d_{0}\|+\varepsilon\leq c_{1}\textup{dist}(x_{0},X^{*})+\varepsilon\leq(1+c_{1})\varepsilon<\frac{\delta}{2}.$
For $k\geq 1$, assume that $x_{i}\in B(x^{*},\frac{\delta}{2})$ and $\textup{dist}(x_{i-1},X^{*})\leq\varepsilon$ for $i=1,...,k$. Then, from Lemma 8 applied to $x_{k-1}$ and $x_{k}$, we obtain
$\|x_{k}-\bar{x}_{k}\|=\textup{dist}(x_{k},X^{*})\leq\eta\textup{dist}(x_{k-1},X^{*})\leq\eta\varepsilon<\varepsilon.$
Moreover, by Lemma 8 and the induction hypothesis for $i\leq k$, we have
$\textup{dist}(x_{i},X^{*})\leq\eta\textup{dist}(x_{i-1},X^{*})\leq\cdots\leq\eta^{i}\textup{dist}(x_{0},X^{*})\leq\eta^{i}\varepsilon<\varepsilon.$
Thus, using Lemma 5,
$\|x_{k+1}-x^{*}\|$ $\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}\|d_{i}\|\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}c_{1}\textup{dist}(x_{i},X^{*})$ (46)
$\leq(1+c_{1})\varepsilon+c_{1}\varepsilon\sum_{i=1}^{\infty}\eta^{i}=\Big{(}1+\frac{c_{1}}{1-\eta}\Big{)}\varepsilon\leq\frac{\delta}{2}.$
Theorem 2.
Suppose that Assumptions 1-3 are valid, and $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$. Let $\{x_{k}\}$ be generated by the LMMSS method with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ for all $k$. If Assumption 4 is verified and $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ is given by (45), then $\{\textup{dist}(x_{k},X^{*})\}$ converges superlinearly to zero. Moreover, the sequence $\{x_{k}\}$ converges to a solution $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$.
Proof.
@@
and since $\textup{dist}(x_{k},X^{*})^{r}\rightarrow 0$ (due to Lemmas 8 and 9) we have that $\{\textup{dist}(x_{k},X^{*})\}$ converges to zero superlinearly.
$\|x_{\ell}-x_{q}\|\leq\sum_{i=q}^{\ell-1}\|d_{i}\|\leq\sum_{i=q}^{\infty}\|d_{i}\|\leq c_{1}\varepsilon\sum_{i=q}^{\infty}\eta^{i},$
implying that $\{x_{k}\}\subset\mathbb{R}^{n}$ is a Cauchy sequence and, therefore, converges. ∎∎
Remark 3.
In particular, when $r=1$ in Assumption 4, $\textup{dist}(x_{k},X^{*})$ converges to zero quadratically.

### v2

@@ 3 Auxiliary Results
Now let us show that $d_{R}\notin\mathcal{N}(L)$. Suppose $d_{R}\in\mathcal{N}(L)$, from (21), we get $\left\lVert Ld_{N}\right\rVert^{2}=-d_{N}^{T}L^{T}Ld_{R}=0,$ implying that $d_{N}\in\mathcal{N}(L)$ and $d_{N}\in\mathcal{N}(J_{k})$, which is a contradiction with condition (5).
By condition (4), there exists $\gamma>0$ such that
$\gamma\left\lVert d_{N}\right\rVert^{2}\leq\left\lVert J_{k}d_{N}\right\rVert^{2}+\left\lVert Ld_{N}\right\rVert^{2},$
which implies
@@
$\left\lVert d_{k}\right\rVert^{2}$ $=\left\lVert d_{N}+d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+2d_{N}^{T}d_{R}+\left\lVert d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}$
@@
$\leq\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}\right\rVert^{2}$
$=\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}-J_{*}^{T}F_{*}\right\rVert^{2}$
@@
Since $J(x_{k})^{T}J(x_{k})$ is continuous, it is clear that for a sufficiently small $\delta>0$, condition (11) is satisfied for all $x_{k}\in B(x^{*},\delta)$ and applying Lemma 2 we have
@@
Lemma 6
Suppose that Assumptions 1 and 2 are valid in $B(x^{*},\delta)$, for some $\delta\in]0,1[$ and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$. If $rank(J_{k})=\ell\geq rank(J_{*})=q\geq 1$, and
(a) $x_{k}\in B(x^{*},\delta)\setminus X^{*}$, Assumption 4 is satisfied with $r\in]0,1]$, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, or,
(b) $x_{k}\in B(x^{*},\bar{\delta})\setminus X^{*}$, Assumption 3 is satisfied with $\sigma<(\sigma_{\min}^{*})^{2}$, where $\sigma_{\min}^{*}$ is the smallest positive singular value of $J(x^{*})$ and
$L_{4}\|x_{k}-\bar{x}_{k}\|+\sigma\leq\lambda_{k}\leq\theta\left(L_{4}\|x_{k}-\bar{x}_{k}\|+\sigma\right),$ (27)
where $\theta>1$ is such that $(\sigma_{\min}^{*})^{2}>\theta\sigma$ and $\bar{\delta}=\min\left\{\delta,\frac{(\sigma_{\min}^{*})^{2}-\theta\sigma}{\theta L_{4}}\right\}$,
then there exists $c_{1}>0$ such that $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$.
Proof
Consider the GSVD for the pair $(J_{k},L)$. In this case where $rank(J_{k})=\ell\geq q=rank(J_{*})\geq 1$ and $rank(L)=p$ with $n-q\leq p$ and $n-\ell\leq p$ (see Remark 2), we observe that:
@@
  • For $n-q\leq i\leq p$, then $(\Gamma_{k})_{ii}=\frac{1}{\sigma_{i,k}^{2}+\lambda_{k}\mu_{i,k}^{2}}<\frac{1}{\sigma_{i,k}^{2}}\leq\frac{\kappa}{(\sigma^{*}_{\text{min}})^{2}},$ where $\sigma^{*}_{\text{min}}$ is the smallest positive singular value of $J(x^{*})$ and $\kappa>1$ (cf. Lemma 2).
Note that, for $x_{k}\in B(x^{*},\delta)$ and due to the continuity of $\mu_{i,k}$, in the case where $n-\ell\leq i<n-q$ we can assume that $\mu_{i,k}>\frac{1}{\sqrt{\kappa}}$ and thus, for $\delta$ sufficiently small, we have $(\Gamma_{k})_{ii}\leq\frac{\kappa}{\lambda_{k}}$.
(a) If $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, for $\delta$ small enough, $\lambda_{k}<\min\{(\sigma_{\min}^{*})^{2}/\kappa,1\}$, and we obtain
$\left\lVert\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]\right\rVert\leq\max{\Big{[}\frac{1}{\lambda_{k}},\frac{\kappa}{\lambda_{k}},\frac{\kappa}{(\sigma^{*}_{\text{min}})^{2}},1\Big{]}}\leq\frac{\kappa}{\lambda_{k}}.$ (28)
From (14) and (28), we get
@@
$=\left\|X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}\Big{(}{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-\right.$
$\qquad\left.-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})+J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\Big{)}\right\|$
$\leq\left\lVert X_{k}\right\rVert^{2}\frac{\kappa}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$
@@
$\left\lVert d_{k}\right\rVert$ $\leq\left\lVert X_{k}\right\rVert^{2}\frac{\kappa}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$ (31)
@@
$\leq\frac{\kappa}{\gamma\lambda_{k}}L_{4}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{2}+\frac{\kappa}{\gamma\lambda_{k}}\left\lVert J_{k}^{T}F(\bar{x}_{k})\right\rVert+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
From Assumptions 2 and 4, the choice $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$ and (9), we obtain from (31) that
$\left\lVert d_{k}\right\rVert$ $\leq\left(\frac{\kappa L_{4}}{\gamma\omega^{r}}+\frac{\kappa C}{\gamma\omega^{r}}+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\right)\left\lVert x_{k}-\bar{x}_{k}\right\rVert,$
which concludes the proof for (a) with $c_{1}=\kappa(L_{4}+C)/\gamma\omega^{r}+\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}/\sqrt{\gamma}$.
(b) Let us denote
$\Theta:=\max{\Big{[}\frac{1}{\lambda_{k}},\frac{\kappa}{\lambda_{k}},\frac{\kappa}{(\sigma^{*}_{\text{min}})^{2}},1\Big{]}},$
Instead of (31), we may write
$\|d_{k}\|\leq\frac{\Theta}{\gamma}L_{4}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{2}+\frac{\Theta}{\gamma}\left\lVert J_{k}^{T}F(\bar{x}_{k})\right\rVert+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
Then, from Assumption 3, we have
$\left\lVert d_{k}\right\rVert$ $\leq\left(\frac{\Theta}{\gamma}(L_{4}\|x_{k}-\bar{x}_{k}\|+\sigma)+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\right)\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
Since, $\|x_{k}-\bar{x}_{k}\|\leq\delta\leq\frac{(\sigma_{\min}^{*})^{2}-\theta\sigma}{\theta L_{4}}$, it follows that $\lambda_{k}\leq(\sigma_{\min}^{*})^{2}$. Then, $\kappa/\lambda_{k}\geq\kappa/(\sigma_{\min}^{*})^{2}$. Clearly, $\kappa/\lambda_{k}>1/\lambda_{k}$ because $\kappa>1$. Hence, $\Theta=\max\{1,\kappa/\lambda_{k}\}$.
Now, assume $\lambda_{k}$ from (27). In case $\Theta=\kappa/\lambda_{k}$, the above inequality becomes
$\left\lVert d_{k}\right\rVert\leq\left(\frac{\kappa}{\gamma}+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\right)\left\lVert x_{k}-\bar{x}_{k}\right\rVert,$
If $\Theta=1$, from the upper bound on $\delta$ in (b) we obtain
$\left\lVert d_{k}\right\rVert\leq\left(\frac{(\sigma_{\min}^{*})^{2}}{\theta\gamma}+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\right)\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
Therefore, we get the proof for (b) with $c_{1}=(1/\gamma)\max\{\kappa,(\sigma_{\min}^{*})^{2}/\theta\}+\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}/\sqrt{\gamma}$. ∎
@@ 4 Local convergence
Now we focus on the local convergence of the “pure” LMMSS iteration, i.e iteration (2)–(3) with $\alpha_{k}=1$. We assume the initial point $x_{0}$ is in a neighborhood of a (possibly nonisolated) stationary point $x^{*}\in X^{*}$.
Lemmas 5 and 6 showed that $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$. As we shall see in this section, such inequality is key for the local convergence analysis under an error bound condition (Assumption 2). As we mentioned in the introduction, the analysis developed in this section is based on the work behling2019local and extends it to handle singular scaling in LM methods.
@@
Lemma 7
Let $x^{*}\in X^{*}$ and $\{x_{k}\}$ be the sequence generated by the LMMSS (with $\alpha_{k}=1$). Suppose Assumptions 1 and 2 holds in $B(x^{*},\delta)$, for some $\delta\in]0,1[$ and that the uniform completeness condition is valid in $B(x^{*},\delta)\setminus X^{*}$. If $x_{k+1},x_{k}\in B(x^{*},\delta)$ and $\left\lVert d_{k}\right\rVert\leq c_{1}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$, then
@@
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq L_{4}\|d_{k}\|^{2}+L_{0}L_{2}\|x_{k}-\bar{x}_{k}\|\|x_{k+1}-\bar{x}_{k}\|+\|J_{k}^{T}F(\bar{x}_{k})\|$
@@ 4.1 Constant Rank
In this section, we assume that $1\leq\text{rank}(J(x^{*}))=q\leq\min\{n,m\}$, and $\text{rank}(J(x))=\text{rank}(J(x^{*}))$ for each $x\in B(x^{*},\delta)$. From Lemma 5, we recall that $\left\|d_{k}\right\|\leq c_{1}\text{dist}(x_{k},X^{*})$, with $c_{1}=\Big{(}\sqrt{\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1}\Big{)}\frac{\kappa}{\lambda_{p}^{*}}L_{3}$.
@@
Lemma 8
Let $x^{*}\in X^{*}$, suppose that Assumptions 1 and 2 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$, and the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$. Assume $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$, that Assumption 3 is verified with $\sigma<\bar{\sigma}:=\omega/(2+c_{1})$, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|$. If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, where
$\varepsilon\leq\frac{\eta\omega-(2+c_{1})\sigma}{L_{6}},$
with $\eta\in\left]\frac{\sigma(2+c_{1})}{\omega},1\right[$ and $L_{6}=L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1}$, then
@@
Proof
Since $J(x)$ has constant rank in $B(x^{*},\delta)$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, and Assumption 3 is satisfied, from (32) and (6), we have
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})||x_{k}-\bar{x}_{k}||^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||$
@@
$+\sigma\|x_{k}-\bar{x}_{k}\|+\sigma\|x_{k+1}-\bar{x}_{k}\|$
$\leq(L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1})\|x_{k}-\bar{x}_{k}\|^{2}+\sigma\|x_{k}-\bar{x}_{k}\|$
$+(1+c_{1})\sigma\|x_{k}-\bar{x}_{k}\|$
$\leq(L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1})\|x_{k}-\bar{x}_{k}\|^{2}+(2+c_{1})\sigma\|x_{k}-\bar{x}_{k}\|$
$\leq\Big{[}(L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1})\varepsilon+(2+c_{1})\sigma\Big{]}\|x_{k}-\bar{x}_{k}\|.$
As $\sigma<\frac{\omega}{2+c_{1}}$, for any $\eta\in(\frac{\sigma(2+c_{1})}{\omega},1)$, we have $\eta\omega-(2+c_{1})\sigma>0$. Denoting $L_{6}:=L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1}$, for $\varepsilon\leq\frac{\eta\omega-(2+c_{1})\sigma}{L_{6}}$, we obtain $\textup{dist}(x_{k+1},X^{*})\leq\eta\textup{dist}(x_{k},X^{*})$. ∎
Lemma 9
(behling2019local, , Lemma 4.3) Suppose that the assumptions of Lemma 8 are satisfied, and consider
$\varepsilon=\min\left\{\frac{\frac{\delta}{2}}{1+\frac{c_{1}}{1-\eta}},\frac{\eta\omega-(2+c_{1})\sigma}{L_{6}}\right\}.$ (37)
Proof
The proof is done by induction. For $k=0$, we have $\textup{dist}(x_{0},X^{*})\leq\|x_{0}-x^{*}\|\leq\varepsilon$, thus
$\|x_{1}-x^{*}\|\leq\|x_{1}-x_{0}\|+\|x_{0}-x^{*}\|\leq\|d_{0}\|+\varepsilon\leq c_{1}\textup{dist}(x_{0},X^{*})+\varepsilon\leq(1+c_{1})\varepsilon<\frac{\delta}{2},$
and by Lemma  8, $\textup{dist}(x_{1},X^{*})\leq\eta\textup{dist}(x_{0},X^{*})\leq\varepsilon$.
From the induction hypothesis for $i\leq k$, $x_{i}\in B(x^{*},\frac{\delta}{2})$ and $\textup{dist}(x_{i-1},X^{*})\leq\varepsilon$, for $i=1,...,k$. Then, by successively applying Lemma 8 to $x_{i-1}$ and $x_{i}$, for $i=1,...,k$, we have
$\textup{dist}(x_{k},X^{*})\leq\eta\textup{dist}(x_{k-1},X^{*})\leq\cdots\leq\eta^{k}\textup{dist}(x_{0},X^{*})\leq\eta^{k}\varepsilon<\varepsilon.$
Finally, using Lemma 5,
$\|x_{k+1}-x^{*}\|$ $\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}\|d_{i}\|\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}c_{1}\textup{dist}(x_{i},X^{*})$
$\leq(1+c_{1})\varepsilon+c_{1}\varepsilon\sum_{i=1}^{\infty}\eta^{i}=\Big{(}1+\frac{c_{1}}{1-\eta}\Big{)}\varepsilon\leq\frac{\delta}{2},$
which completes the proof. ∎
Theorem 2
Let $x^{*}\in X^{*}$, suppose that Assumptions 1, 2 and 3 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$ and $\bar{\sigma}=\omega/(2+c_{1})$, $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$, and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$. Let $\{x_{k}\}$ be generated by the LMMSS method with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ for all $k$, and $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ is given by (37). Then, $\{\textup{dist}(x_{k},X^{*})\}$ converges linearly to zero. Moreover, the sequence $\{x_{k}\}$ converges to some $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$.
Proof
The linearly convergence of $\textup{dist}(x_{k},X^{*})$ to zero follows directly from Lemma 8 and Lemma 9.
$\|x_{\ell}-x_{q}\|\leq\sum_{i=q}^{\ell-1}\|d_{i}\|\leq\sum_{i=q}^{\infty}\|d_{i}\|\leq c_{1}\varepsilon\sum_{i=q}^{\infty}\eta^{i}\leq\frac{c_{1}\varepsilon}{1-\eta},$
implying that $\{x_{k}\}\subset\mathbb{R}^{n}$ is a Cauchy sequence and, therefore, converges. ∎
Now we demonstrate that under Assumption 4, which is stronger than Assumption 3, the local convergence can be superlinear. In the lemmata and theorems that follows we shall suppress algebraic manipulations that are similar to the results above.
Lemma 10
Let $x^{*}\in X^{*}$, suppose that Assumptions 1 and 2 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$, $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$, and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$. Additionally, assume that Assumption 4 is verified, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|$. If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, with
$\varepsilon\leq\left(\frac{\eta\omega}{\tilde{C}}\right)^{1/r},$
with $\eta\in]0,1[$, $\tilde{C}=L_{6}+(1+(1+c_{1})^{1+r})C$ and $L_{6}$ as in Lemma 8, then
$\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*}).$
Proof
Similarly to the proof of Lemma 8, we have
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{2}$ (38)
@@
$\textup{dist}(x_{k+1},X^{*})\leq\eta\textup{dist}(x_{k},X^{*}),$
which concludes the proof. ∎
Lemma 11
@@
Proof
Same as that of Lemma 9. ∎
Theorem 3
Let $x^{*}\in X^{*}$, suppose that Assumptions 1, 2 and 4 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$, $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$, and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$. Let $\{x_{k}\}$ be generated by the LMMSS method with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ for all $k$, and $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ is given by (39). Then, $\{\textup{dist}(x_{k},X^{*})\}$ converges superlinearly to zero. Moreover, the sequence $\{x_{k}\}$ converges to some $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$.
Proof
@@
and since $\textup{dist}(x_{k},X^{*})^{r}\rightarrow 0$ (due to Lemmas 10 and 11) we have that $\{\textup{dist}(x_{k},X^{*})\}$ converges to zero superlinearly.
Proof of convergence of the whole sequence $\{x_{k}\}$ to some $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$ follows the same lines as in the proof of Theorem 2. ∎